### v2

# Improving Probabilistic Diffusion Models With Optimal Covariance Matching

Abstract
The probabilistic diffusion model has become highly effective across various domains. Typically, sampling from a diffusion model involves using a denoising distribution characterized by a Gaussian with a learned mean and either fixed or learned covariances. In this paper, we leverage the recently proposed covariance moment matching technique and introduce a novel method for learning the diagonal covariances. Unlike traditional data-driven covariance approximation approaches, our method involves directly regressing the optimal analytic covariance using a new, unbiased objective named Optimal Covariance Matching (OCM). This approach can significantly reduce the approximation error in covariance prediction. We demonstrate how our method can substantially enhance the sampling efficiency, recall rate and likelihood of both diffusion models and latent diffusion models.
1 Introduction
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song & Ermon, 2019) have achieved remarkable success in modeling complex real-world data across various domains (Rombach et al., 2022; Li et al., 2022; Poole et al., 2022; Ho et al., 2022; Hoogeboom et al., 2022; Liu et al., 2023). A conventional diffusion model operates in two stages: a forward noising process, indexed by $t\in[1,T]$, which progressively corrupts the data distribution into a Gaussian distribution via Gaussian convolution, and a reverse denoising process, which generates images by gradually transforming Gaussian noise back into coherent data samples. In traditional diffusion models, the generation process typically predicts only the mean of the denoising distribution while using a fixed, pre-defined variance (Ho et al., 2020). This approach often requires a very large number of steps, $T$, to produce high-quality, diverse samples or to achieve reasonable model likelihoods, leading to inefficiencies during inference.
To address this inefficiency, several works have proposed estimating the posterior covariance instead of relying on pre-defined variance values. For instance, Bao et al. (2022b) introduces an analytical form of isotropic, state-independent covariance that can be estimated from the data. This analytical solution achieves an optimal form under the constraints of isotropy and state independence. A more flexible approach involves learning a state-dependent diagonal covariance. Nichol & Dhariwal (2021) propose learning this form by optimizing the variational lower bound (VLB). Similarly, Bao et al. (2022a) explore learning a state-dependent diagonal covariance directly from data, also examining its analytical form. These methods have demonstrated improved image quality with fewer denoising steps. Learning the covariance through VLB optimization (Nichol & Dhariwal, 2021) has become a widely adopted strategy in state-of-the-art image and video generative models.
Building on this line of research, the goal of our paper is to develop an improved denoising covariance strategy that enhances both the generation quality and likelihood evaluation while reducing the number of total time steps. Recently, Zhang et al. (2024b) derived the optimal state-dependent full covariance for the denoising distribution. While this method offers greater flexibility than state-dependent diagonal covariance, it requires $O(D^{2})$ storage for the Hessian matrix and $O(D)$ network evaluations per denoising step, where $D$ is the data dimension. This makes it impractical for high-dimensional applications, such as image generation. To address this limitation, we propose a novel, unbiased covariance matching objective that enables a neural network to match the diagonal of the optimal state-dependent diagonal covariance. Unlike previous methods (Bao et al., 2022a; Nichol & Dhariwal, 2021), which learn the diagonal covariance directly from the data, our approach estimates the diagonal covariance from the learned score function. We show that this method significantly reduces covariance estimation errors compared to existing techniques. Moreover, we demonstrate that our approach can be applied to both Markovian (DDPM) and non-Markovian (DDIM) diffusion models, as well as latent diffusion models. This results in improvements in generation quality, recall, and likelihood evaluation, while also reducing the number of function evaluations (NFEs).
2 Background of Probabilistic Diffusion Models
We first introduce two classes of diffusion models that will be explored further in our paper.
2.1 Markovian Diffusion Model: DDPM
Let $q(x_{0})$ be the true data distribution, denoising diffusion probabilistic models (DDPM) (Ho et al., 2020) constructs the following Markovian diffusion process
$q(x_{0:T})=q(x_{0})\prod_{t=1}^{T}q(x_{t}|x_{t-1}),$ (1)
where $q(x_{t}|x_{t-1})=\mathcal{N}(\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$ and $10{<}\beta_{1}{<}\cdots{<}\beta_{T}{<}$ 1 is the pre-defined noise schedule. We can also derive the following skip-step noising distribution in a closed form
$q(x_{t}|x_{0})=\mathcal{N}(\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})I),$ (2)
where $\bar{\alpha}_{t}\equiv\prod_{s=1}^{t}(1-\beta_{s})$. When $T$ is sufficiently large, we have the marginal distribution $q(x_{T})\rightarrow\mathcal{N}(0,I)$. The generation process utilizes an initial distribution $q(x_{T})$ and the denoising distribution $q(x_{t-1}|x_{t})$. We assume for a large $T$ and variance preversing schedule discussed in Equation 1, we can approximate $q(x_{T})\approx p(x_{T})=\mathcal{N}(0,I)$. The true $q(x_{t-1}|x_{t})$ is intractable and is approximated as a variational distribution $p_{\theta}(x_{t-1}|x_{t})$ within a Gaussian family, which defines the reverse denoising process as
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1}|\mu_{t-1}(x_{t};\theta),\Sigma_{t-1}(x_{t};\theta)).$ (3)
Using the Tweedie’s Lemma (Efron, 2011; Robbins, 1992), we can obtain the score representation of the $\mu_{t-1}(x_{t};\theta)$
$\mu_{t-1}(x_{t};\theta)=(x_{t}+\beta_{t}\nabla_{x_{t}}\log p_{\theta}(x_{t}))/\sqrt{1-\beta_{t}},$ (4)
where the approximated score function $\nabla_{x_{t}}\log p_{\theta}(x_{t})\approx\nabla_{x_{t}}\log q(x_{t})$ can be learned by the denoising score matching (DSM) (Vincent, 2011; Song & Ermon, 2019).
For the covariance $\Sigma_{t-1}(x_{t};\theta)$, two heuristic choices were proposed in the original DDPM paper: 1. $\Sigma_{t-1}(x_{t};\theta)=\beta_{t}$, which is equal to the variance of $q(x_{t}|x_{t-1})$ and 2. $\Sigma_{t-1}(x_{t};\theta)=\tilde{\beta}_{t}$ where $\tilde{\beta}_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}$ is the variance of $q(x_{t-1}|x_{t},x_{0})$. Although heuristic, Ho et al. (2020) show that when $T$ is large, both options yield similar generation quality.
2.2 Non-Markovian Diffusion Model: DDIM
In additional to the Markovian diffusion process, denoising diffusion implicit models (DDIM) (Song et al., 2020) only defines the condition $q(x_{t}|x_{0})=\mathcal{N}(\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})I)$ and let
[…]
where $q(x_{t-1}|x_{t},x_{0})=\mathcal{N}(\mu_{t-1},\sigma^{2}_{t-1})$ with
$\mu_{t{-}1}=\sqrt{\bar{\alpha}_{t{-}1}}x_{0}+\frac{x_{t}{-}\sqrt{\bar{\alpha}_{t}}x_{0}}{\sqrt{1-\bar{\alpha}_{t}}}\sqrt{1{-}\bar{\alpha}_{t{-}1}{-}\sigma_{t{-}1}^{2}}.$ (6)
When $\sigma_{t-1}=\sqrt{(1-\bar{\alpha}_{t-1})/(1-\bar{\alpha}_{t})\beta}_{t}$, the diffusion process becomes Markovian and equivalent to DDPM. Specifically, DDIM chooses $\sigma\rightarrow 0$, which implicitly defines a non-Markovian diffusion process. In the original paper (Song et al., 2020), the $q(x_{0}|x_{t})$ is heuristically chosen as a delta function $p_{\theta}(x_{0}|x_{t})=\delta(x_{0}-\mu_{0}(x_{t};\theta))$ where
$\mu_{0}(x_{t};\theta)=(x_{t}+(1-\bar{\alpha}_{t})\nabla_{x_{t}}\log p_{\theta}(x_{t}))/\sqrt{\bar{\alpha}_{t}}.$ (7)
In both DDPM and DDIM, the covariance of $p_{\theta}(x_{t-1}|x_{t})$ or $p_{\theta}(x_{0}|x_{t})$ are chosen based on heuristics. Nichol & Dhariwal (2021) have shown that the choice of covariance makes a big impact when $T$ is small. Therefore, for the purpose of accelerating the diffusion sampling, our paper will focus on how to improve the covariance estimation quality in these cases. We will first introduce our method in the next section and then compare it with other methods in Section 4.
3 Diffusion Models with Optimal Covariance Matching
Recently, Zhang et al. (2024b) introduce the optimal covariance form of the denoising distribution $q(x|\tilde{x})\propto q(\tilde{x}|x)q(x)$ for the Gaussian convolution $q(\tilde{x}|x)=\mathcal{N}(x,\sigma^{2}I)$, which can be seen as a high-dimensional extension of the second-order Tweedie’s Lemma (Efron, 2011; Robbins, 1992). We further extend the formula to scaled Gaussian convolutions in the following theorem.
Theorem 1 (Generalized Analytical Covariance Identity).
Given a joint distribution $q(\tilde{x},x)=q(\tilde{x}|x)q(x)$ with $q(\tilde{x}|x)=\mathcal{N}(\alpha x,\sigma^{2}I)$, then the covariance of the true posterior $q(x|\tilde{x})\propto q(x)q(\tilde{x}|x)$, which is defined as $\Sigma(\tilde{x})=\mathbb{E}_{q(x|\tilde{x})}[x^{2}]-\mathbb{E}_{q(x|\tilde{x})}[x]^{2}$, has a closed form:
$\Sigma(\tilde{x})=\left(\sigma^{4}\nabla_{\tilde{x}}^{2}\log q(\tilde{x})+\sigma^{2}I\right)/\alpha^{2}.$ (8)
See Appendix A.1 for a proof. This covariance form can also be shown as the optimal covariance form under the KL divergences (Bao et al., 2022b; Zhang et al., 2024b). We can see that the exact covariance in this case only depends on the score function, which indicates the exact covariance can be derived from the score function. In general, the score function already contains all the information of the denoising distribution $q(x|\tilde{x})$, see Zhang et al. (2024b) for further discussion. We here only consider the covariance for simplicity.
Table 1: Comparison of different covariance choices on CIFAR10 (CS). We use Rademacher estimator ($M=100$) for the diagonal covariance approximation in OCM-DDPM.
FiD $\downarrow$ NLL $\downarrow$
# timesteps 5 10 15 20 5 10 15 20
DDPM, $\tilde{\beta}$ 58.28 34.76 24.02 19.00 203.29 74.95 44.94 32.20
DDPM, ${\beta}$ 254.07 205.31 149.67 109.81 7.33 6.51 6.06 5.77
OCM-DDPM 38.88 21.60 13.35 9.75 6.82 4.98 4.62 4.43
In the case of the diffusion model, we use the learned score function as a plug-in approximation in Equation 8. Although the optimal covariance can be directly calculated from the learned score function, but it requires calculating the Hessian matrix, which is the Jacobian of the score function. This requires $O(D^{2})$ storage and $D$ number of network evaluation (Martens et al., 2012) for each denoising step at the time $t$. Zhang et al. (2024b) propose to use the following consistent diagonal approximation (Bekas et al., 2007) to remove the $O(D^{2})$ storage requirement
$\mathrm{diag}(H(\tilde{x}))\approx 1/M\sum\nolimits_{m=1}^{M}v_{m}\odot H(\tilde{x})v_{m},$ (9)
where $H(\tilde{x})\equiv\nabla_{\tilde{x}}^{2}\log q(\tilde{x})$ and $v_{m}{\sim}p(v)$ is a Rademacher random variable (Hutchinson, 1990) with entries ±$\pm 1$ and $\odot$ denotes the element-wise product. In Table 1, we compare the generation quality of DDPM using different covariance choices. The results demonstrate that generation quality improves significantly when using the Rademacher estimator to estimate the optimal covariance, compared to the heuristic choices of $\beta$ and $\tilde{\beta}$. However, as shown in Figure 1, achieving a desirable approximation on the CIFAR10 dataset necessitates $M\geq 100$ Rademacher samples. Each calculation of $v_{m}\odot H(\tilde{x})v_{m}$ requires two network passes, resulting in a total of $2M$ network evaluations. This significantly slows down the generation speed, making it impractical for diffusion models. Inspired by Nichol & Dhariwal (2021); Bao et al. (2022a) and also the amortization technique used in variational inference (Kingma & Welling, 2013; Dayan et al., 1995), we propose to use a network to match the diagonal Hessian, which only requires one network pass to predict the diagonal Hessian in the generation time and can be done in parallel with the score/mean predictions with no extra time cost. In the next section, we introduce a novel unbiased objective to learn the diagonal Hessian from the learned score, which improves the covariance estimation accuracy and leads to better generation quality and higher likelihood estimations.
3.1 Unbiased Optimal Covariance Matching
Data
$K=1$
$K=10$
$K=100$
$K=1000$
Figure 1: Diagonal covariance estimation visualisation with different Rademacher sample numbers.
To train a network $h_{\phi}(\tilde{x})$ to match the Hessian diagonal $\text{diag}(H(\tilde{x}))$, a straightforward solution is to directly regress Equation 9 for all the noisy data
$\min_{\phi}\mathbb{E}_{q(\tilde{x})}||h_{\phi}(\tilde{x})\!-\!\frac{1}{M}\sum_{m=1}^{M}v_{m}\!\odot\!H(\tilde{x})v_{m}||_{2}^{2},$ (10)
where $v_{m}\sim p(v)$. Although this objective is consistent when $M\rightarrow\infty$, it will introduce additional bias when $M$ is small. To avoid the bias, we propose the following unbiased optimal covariance matching (OCM) objective
$L_{\text{ocm}}(\phi)=\mathbb{E}_{q(\tilde{x})p(v)}||h_{\phi}(\tilde{x})-v\odot H(\tilde{x})v||_{2}^{2},$ (11)
which does not include an expectation within the non-linear L2 norm. The following theorem shows the validity of the proposed OCM objective.
Theorem 2 (Validity of the OCM objetcive).
The objective in Equation 11 upper bounded the grounded objective. Moreover, it attains optimal when $h_{\phi}(\tilde{x})=\text{diag}(H(\tilde{x}))$ for all $\tilde{x}\sim q(\tilde{x})$.
See the Section A.2 for a proof. The integration over $v$ in Equation 11 can be unbiasedly approximated by the Monte-Carlo integration given $M$ Rademacher samples $v_{m}\sim p(v)$. In practice, we found $M=1$ works well, which also shows the training efficiency of the proposed OCM objective. The learned $h_{\phi}$ can form the covariance approximation
$\Sigma(\tilde{x};\phi)=(\sigma^{4}h_{\phi}(\tilde{x})+\sigma^{2}I)/\alpha^{2}.$ (12)
We then discuss how to apply the learned covariance to diffusion models.
Algorithm 1 OCM-DDPM
1:  Compute $\mu_{t^{\prime}}(x_{t})$ with Equation 15;
2:  Compute $\Sigma_{t^{\prime}}(x_{t})$ with Equation 16;
3:  Sample $\epsilon\!\sim\!\mathcal{N}(0,I)$;
4:  Let $x_{t^{\prime}}\leftarrow\mu_{t^{\prime}}(x_{t})+\Sigma_{t^{\prime}}(x_{t})^{1/2}\epsilon$.
Algorithm 2 OCM-DDIM
1:  Compute $\mu_{0}(x_{t})$ with Equation 18;
2:  Compute $\Sigma_{0}(x_{t})$ with Equation 19;
3:  Sample $x_{0}\!\sim\!\mathcal{N}(\mu_{0}(x_{t}),\!\Sigma_{0}(x_{t}))$;
4:  Let $x_{t^{\prime}}\leftarrow\sqrt{\bar{\alpha}}_{t^{\prime}}x_{0}\!+\sqrt{1-\bar{\alpha}_{t^{\prime}}}\frac{x_{t}-\sqrt{\bar{\alpha}_{t}}x_{0}}{\sqrt{1-\bar{\alpha}_{t}}}$.
Figure 2: The sampling procedure OCM-DDPM and OCM-DDIM. We only demonstrate one step from $t\rightarrow t^{\prime}$ for simplicity.
3.2 Diffusion Models Applications
Given that we have access to a learned score function $\nabla_{x_{t}}\log p_{\theta}(x_{t})$, which can be obtained from any pre-trained diffusion model, we denote the Jacobian of the score by $H_{t}(x_{t})$. Assuming $M=1$ in the OCM training objective, the covariance learning objective for diffusion models can then be expressed as follows
$\min_{\phi}\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}_{q(x_{t},x_{0})p(v)}\lVert h_{\phi}(x_{t})-v{\odot}H_{t}(x_{t})v\rVert_{2}^{2},$ (13)
where $v\sim p(v)$ and $h_{\phi}(x_{t})$ is a network that conditioned on the state $x_{t}$ and time $t$. After training this objective, the learned $h_{\phi}(x_{t})$ can be used to form the diagonal Hessian approximation $h_{\phi}(\tilde{x})\approx\mathrm{diag}(H(\tilde{x}))$ which further forms our approximation of covariance. We then derive its use cases in skip-step DDPM and DDIM for Diffusion acceleration.
Skip-Step DDPM: For the general skip-step DDPM with denoising distribution $q(x_{t^{\prime}}|x_{t})$ with $t^{\prime}<t$. When $t^{\prime}=t-1$, this becomes the classic one-step DDPM. We further denote $\bar{\alpha}_{t^{\prime}:t}=\prod_{s=t^{\prime}}^{t}\alpha_{s}$, and thus $\bar{\alpha}_{0:t}=\bar{\alpha}_{t}$, $\bar{\alpha}_{t^{\prime}:t}=\bar{\alpha}_{t}/\bar{\alpha}_{t^{\prime}}$. We can write the forward process as
$q(x_{t}|x_{t^{\prime}})=\mathcal{N}(x_{t}|\sqrt{\bar{\alpha}_{t^{\prime}:t}}x_{t^{\prime}},(1-\bar{\alpha}_{t^{\prime}:t})I).$ (14)
The corresponding Gaussian denoising distribution $p_{\theta,\phi}(x_{t^{\prime}}|x_{t})=\mathcal{N}(\mu_{t^{\prime}}(x_{t};\theta),\Sigma_{t^{\prime}}(x_{t};\phi))$ has the following mean and covariance functions:
$\mu_{t^{\prime}}(x_{t};\theta)$ $\!=\!(x_{t}+(1-\bar{\alpha}_{t^{\prime}:t})\nabla_{x_{t}}\log p_{\theta}(x_{t}))/\sqrt{\bar{\alpha}_{t^{\prime}:t}},$ (15)
$\Sigma_{t^{\prime}}(x_{t};\phi)$ $\!=\!({(1\!-\!\bar{\alpha}_{t^{\prime}:t})}^{2}h_{\phi}(x_{t})\!+\!(1\!-\!\bar{\alpha}_{t^{\prime}:t})I)/\bar{\alpha}_{t^{\prime}:t}.$ (16)
The skip-step denoising sample $x_{t^{\prime}}\sim p_{\theta,\phi}(x_{t}^{\prime}|x_{t})$ can be obtained by $x_{t^{\prime}}=\mu_{t^{\prime}}(x_{t};\theta)+\epsilon\Sigma^{1/2}_{t^{\prime}}(x_{t};\phi)$.
Skip-Step DDIM: Similarly, we give the skip-step formulation of DDIM, where we set the $\sigma_{t-1}=0$ in Equation 6 which is also used in the original paper Song et al. (2020). We can use the approximated covariance of $p_{\theta,\phi}(x_{0}|x_{t})$ to replace the delta function used in the vanilla DDIM, which gives the skip-steps DDIM sample
$x_{t^{\prime}}$ $=\sqrt{\bar{\alpha}}_{t^{\prime}}x_{0}+\sqrt{1-\bar{\alpha}_{t^{\prime}}}\cdot\frac{x_{t}-\sqrt{\bar{\alpha}_{t}}x_{0}}{\sqrt{1-\bar{\alpha}_{t}}},$ (17)
where $x_{0}\sim p_{\theta,\phi}(x_{0}|x_{t})=\mathcal{N}(\mu_{0}(x_{t};\theta),\Sigma_{0}(x_{t};\phi))$ and
$\mu_{0}(x_{t};\theta)$ $=(x_{t}+(1-\bar{\alpha}_{t})\nabla_{x_{t}}\log p_{\theta}(x_{t}))/\sqrt{\bar{\alpha}_{t}},$ (18)
$\Sigma_{0}(x_{t};\phi)$ $=({(1-\bar{\alpha}_{t})}^{2}h_{\phi}(x_{t},t)+(1-\bar{\alpha}_{t})I)/\bar{\alpha}_{t}.$ (19)
The sample $x_{0}\sim p_{\theta,\phi}(x_{0}|x_{t})$ can be obtained by $x_{0}=\mu_{0}(x_{t};\theta)+\epsilon\Sigma^{1/2}_{0}(x_{t};\phi)$. In the next section, we give an overview of other covariance estimation methods and compare their performance in practical examples.
Table 2: Overview of different covariance estimation methods. Methods are ranked by increasing modeling capability from top to bottom. We also include the intuition of the methods and how many additional network passes are required for estimating the covariance.
Modeling Capability
Covariance Type +#Passes Intuition
$x_{t}$-independent Isotropic $\beta$ (Ho et al., 2020) 0 Cov. of $q(x_{t}|x_{t-1})$
$x_{t}$-independent Isotropic $\tilde{\beta}$ (Ho et al., 2020) 0 Cov. of $q(x_{t}|x_{t-1},x_{0})$
$x_{t}$-independent Isotropic Estimation (Bao et al., 2022b) 0 Estimate from data
$x_{t}$-dependent Diagonal VLB (Nichol & Dhariwal, 2021) 1 Learn from data
$x_{t}$-dependent Diagonal NS (Bao et al., 2022a) 1 Learn from data
$x_{t}$-dependent Diagonal OCM (Ours) 1 Learn from score
$x_{t}$-dependent Diagonal Estimation (Zhang et al., 2024b) $2M$ Estimate from score
$x_{t}$-dependent Diagonal Analytic (Zhang et al., 2024b) $D$ Calculate from score
$x_{t}$-dependent Full Analytic (Zhang et al., 2024b) $D$ Calculate from score
3.3 Details of Training and Inference
Our model comprises two components: a score prediction network $s_{\theta}$ and a diagonal Hessian prediction network $h_{\phi}$. In line with Bao et al. (2022a), we parameterize the score prediction network using a pretrained diffusion model, and the Hessian prediction network is parameterized by sharing parameters as follows:
$s_{\theta}(x_{t})$ $=\mathrm{NN}_{1}(\mathrm{BaseNet}(x_{t},t;\theta_{1});\theta_{2})$ (20)
$h_{\phi}(x_{t})$ $=\mathrm{NN}_{2}(\mathrm{BaseNet}(x_{t},t;\theta_{1});\phi)$ (21)
where $\mathrm{BaseNet}$ represents the commonly used architecture in diffusion models, such as UNet and DiT (Ho et al., 2020; Peebles & Xie, 2023). This parameterization approach only requires an additional small neural network, $\mathrm{NN}_{2}$, resulting in negligible extra computational and memory costs compared to the original diffusion models (see Section B.1 for more details). In our experiment, we fix the parameter $\theta=\{\theta_{1},\theta_{2}\}$ and train the Hessian prediction network exclusively with the proposed OCM objective. After training, samples can be generated using the algorithms summarized in Figure 2.
4 Related Covariance Estimation Methods
We then discuss different choices of $\Sigma_{t^{\prime}}(x_{t})$ used in the diffusion model literature, see also Table 2 for an overview. For brevity, we mainly focus on DDPM here, in which the mean $\mu_{t^{\prime}}(x_{t})$ can be computed as in Equation 15.
1. $x_{t}$-independent isotropic covariance: $\beta$/$\tilde{\beta}$-DDPM (Ho et al., 2020).
The $\beta$-DDPM uses the variance of $p(x_{t^{\prime}}|x_{t})$, which is $\Sigma_{t^{\prime}}(x_{t})=(1-\bar{\alpha}_{t}/\bar{\alpha}_{t^{\prime}})I$, when $t^{\prime}=t-1$, we have $\Sigma_{t-1}(x_{t})=\beta_{t}$. The $\tilde{\beta}$-DDPM uses the covariance of $p(x_{t^{\prime}}|x_{0},x_{t})$, which is $\frac{(1-\bar{\alpha}_{t^{\prime}})}{(1-\bar{\alpha}_{t})}(1-\bar{\alpha}_{t^{\prime}:t})$.
2. $x_{t}$-independent isotropic covariance: A-DDPM (Bao et al., 2022b).
A-DDPM assumes a state-independent isotropic covariance $\Sigma_{t^{\prime}}(x_{t})=\sigma_{t^{\prime}}^{2}I$ with the following analytic form of
$\sigma_{t^{\prime}}^{2}=\frac{1\!-\!\bar{\alpha}_{t^{\prime}:t}}{\bar{\alpha}_{t^{\prime}:t}}\!-\!\frac{(1\!-\!\bar{\alpha}_{t^{\prime}:t})^{2}}{d\bar{\alpha}_{t^{\prime}:t}}\mathbb{E}_{q(x_{t})}\left[\left\lVert\nabla_{\!x_{t}}\!\log p_{\theta}(x_{t})\right\rVert_{2}^{2}\right],$
where the integration of $q(x_{t})$ requires a Monte Carlo estimation before conducting generation. This variance choice is optimal under the KL divergence within a constrained isotropic state-independent posterior family.
3. $x_{t}$-dependent diagonal covariance: I-DDPM (Nichol & Dhariwal, 2021).
In I-DDPM, the diagonal covariance matrix is modelled as the interpolation between $\beta_{t}$ and $\tilde{\beta_{t}}$
$\Sigma_{t-1}(x_{t};\psi)=\exp(v_{\psi}(x_{t})\log\beta_{t}+(1-v_{\psi}(x_{t}))\log\tilde{\beta}_{t}),$ (22)
where $v_{\psi}$ is parametrized via a neural network. The covariance is learned with the variational lower bound (VLB). In the optimal training, the covariance learned by VLB will recover the true covariance in (8). Notably, Nichol & Dhariwal (2021) heuristically obtain the covariance of skip sampling $\Sigma_{t^{\prime}}(x_{t};\psi)$ by rescaling $\beta_{t}$ and $\tilde{\beta_{t}}$ accordingly: $\beta_{t}\rightarrow 1-\bar{\alpha}_{t^{\prime}:t},\tilde{\beta}_{t}\rightarrow\frac{(1-\bar{\alpha}_{t^{\prime}})}{(1-\bar{\alpha}_{t})}(1-\bar{\alpha}_{t^{\prime}:t})$. However, when $t^{\prime}=0$, $\Sigma_{0}(x_{t};\psi)$ is ill-defined; thus, iDDPM is inapplicable within the DDIM framework.
4. $x_{t}$-dependent diagonal covariance: SN-DDPM (Bao et al., 2022a).
SN-DDPM learns the covariance by training a neural network $g_{\psi}$ to estimate the second moment of the noise $\epsilon_{t}=(x_{t}-\sqrt{\bar{\alpha}_{t}}x_{0})/\sqrt{1-\bar{\alpha}_{t}}$:
$\min_{\psi}\mathbb{E}_{t,q(x_{0},x_{t})}\lVert\epsilon^{2}_{t}-g_{\psi}(x_{t})\rVert_{2}^{2}.$ (23)
After training, the covariance $\Sigma_{t^{\prime}}(x_{t};\psi)$ can be estimated via
$\!\!\!\!\Sigma_{t^{\prime}}(x_{t};\psi)\!=\!\frac{(1\!-\!\bar{\alpha}_{t^{\prime}})}{(1-\bar{\alpha}_{t})}\bar{\beta}_{t^{\prime}:t}I\!+\!\frac{\beta_{t}^{2}}{\alpha_{t}}\!\left(\!\frac{g_{\psi}(x_{t})}{1\!-\!\bar{\alpha}_{t}}\!-\!\nabla_{x_{t}}\log\!p_{\theta}(x_{t})^{2}\!\right)\!.$ (24)
where $\bar{\beta}_{t^{\prime}:t}=1-\bar{\alpha}_{t^{\prime}:t}$. In optimal, $\Sigma_{t^{\prime}}(x_{t};\psi)$ in Equation 24 will recover the true covariance in Equation 8. We demonstrate the equivalence between COM-DDPM and SN-DDPM in Section A.3. However, due to the appearance of the quadratic noise term in Equation 23, SN-DDPM tends to amplify the estimation error as $t\rightarrow 0$, leading to suboptimal solutions. To mitigate this issue, Bao et al. (2022a) also propose NPR-DDPM, which models the noise prediction residual instead. We recommend referring to their paper for detailed explanations.
Notably, almost all these methods can be applied within the DDIM framework by setting $t^{\prime}=0$. Specifically, $p(x_{t^{\prime}}|x_{t})$ can be sampled using Equation 17 with $x_{0}\sim\mathcal{N}(\mu_{0}(x_{t}),\Sigma_{0}(x_{t}))$, where $\mu_{0}(x_{t})$ is the same as in (18), but $\Sigma_{0}(x_{t})$ differs for various methods as discussed previously.
(a) Cov Error with True Score
(b) Cov Error with Learned Score
(c) DDPM MMD v.s. Steps
(d) DDIM MMD v.s. Steps
Figure 3: Comparisons of different covariance estimation methods based on estimation error and sample generation quality. Figures (a) and (b) show the mean square error of the estimated diagonal covariance under the assumptions: (a) access to the true score, and (b) learned score of the data distribution at various noise levels. Figures (c) and (d) present the MMD evaluation against the total sampling steps in the DDPM (c) and DDIM (d) settings. We can find the proposed OCM method can achieve the lowest estimation error and consistently outperform other methods when fewer generation steps are applied.
5 Experimental Results
To support our theoretical discussion, we first evaluate the performance of optimal covariance matching by training diffusion probabilistic models on 2D toy examples. We then demonstrate its effectiveness in enhancing image modelling in both pixel and latent spaces, focusing on the comparison between optimal covariance matching and other covariance estimation methods, and showing that the proposed approach has the potential to scale to large image generation tasks.
5.1 Toy Demonstration
We first demonstrate the effectiveness of our method by considering the data distribution as a two-dimensional mixture of nine Gaussians (MoG) with means located at $[-3,0,3]\otimes[-3,0,3]$ and a standard deviation of $\sigma=0.1$, where $\otimes$ denotes Cartesian product. In this case, both the true score $\nabla_{x_{t}}\log p(x_{t})$ and the optimal covariance $\Sigma_{t^{\prime}}(x_{t})$ are available, allowing us to compare the covariance estimation error. All methods, including I-DDPM, A-DDPM, SN-DDPM, and our OCM-DDPM, assume accurate score estimation. We first learn the covariance using the true scores. Specifically, we train the covariance networks for these methods over 50,000 iterations with a learning rate of $0.001$ using an Adam optimizer (Kingma & Ba, 2014). Figure 3(a) shows the $L_{2}$ error of the estimated diagonal covariance $\Sigma_{t{-}1}(x_{t})$, and our method, OCM-DDPM, consistently achieves the lowest error compared to the other methods. In practice, the true score is not accessible. Therefore, we also conduct a comparison using a score function learned with DSM. We then apply different covariance estimation methods under the same settings. In Figure 3(b), we plot the same $L_{2}$ error for the learned score setting and find that our method achieves the lowest error for all $t$ values.
We further use the learned score and the covariance estimated by different methods to conduct DDPM and DDIM sampling for this MoG problem. For evaluation, we employ the Maximum Mean Discrepancy (MMD) (Gretton et al., 2012), which utilizes five kernels with bandwidths $\{2^{-2},2^{-1},2^{0},2^{1},2^{2}\}$. In Figures 3(c) and 3(d), we show the MMD comparison among different methods. Specifically, we choose to conduct different diffusion steps with the skip-step scheme as discussed in Section 3.2. We find that our method consistently outperforms the others when the total time step is small, demonstrating the importance of accurate covariance estimation in the context of diffusion acceleration.
5.2 Image Modelling with Diffusion Models
Following the experimental setting in Bao et al. (2022a), we then evaluate our method across varying pre-trained score networks provided by previous works (Ho et al., 2020; Nichol & Dhariwal, 2021; Song et al., 2020; Bao et al., 2022b; Peebles & Xie, 2023). In this experiment, we mainly focus on four datasets: CIFAR10 (Krizhevsky et al., 2009) with the linear schedule (LS) of $\beta_{t}$ (Ho et al., 2020) and the cosine schedule (CS) of $\beta_{t}$ (Nichol & Dhariwal, 2021); CelebA (Liu et al., 2015); LSUN Bedroom (Yu et al., 2015). The details of the experimental setting can be found in Appendix B.
Improving Likelihood Results
We first evaluate the negative log-likelihood (NLL) of the diffusion models by calculating the negative evidence lower bound (ELBO): $-\log q(x_{0})\leq\mathbb{E}_{q}\log\frac{p(x_{0:T})}{q(x_{1:T}|x_{0})}\equiv-L_{\mathrm{elbo}}(x_{0})$ with the same mean yet varying covariances. As per Bao et al. (2022b), we report results only for the DDPM forward process, as $-L_{\mathrm{elbo}}=\infty$ under the DDIM forward process. As shown in Table 3, our OCM-DDPM demonstrates the best or second-best performance in most scenarios, highlighting the effectiveness of our covariance estimation approach. We also note that SN-DDPM performs poorly on likelihood results in small-scale datasets like CIFAR10 and CELEBA, likely due to the amplified error of the quadratic term in Equation 23. Although NPR-DDPM achieves slightly better performance in certain scenarios, it falls short in terms of sample quality, as measured by FID, shown in Table 5. In Section C.1, we also compare our method to Improved DDPM (Nichol & Dhariwal, 2021) in terms of likelihood on ImageNet 64x64. The results show that our method achieves better likelihood estimation with a small number of timesteps, while achieving comparable results with full timesteps.
Improving Sample Quality
Next, we compare sample quality quantitatively using the FID score, with results reported for both the DDPM and DDIM forward processes. As shown in Table 5, our OCM-DPM consistently achieves either the best or second-best results across most settings. The qualitative results of the samples generated by our method can be found in Section C.4.
Notably, while SN-DDPM often demonstrates slightly better FID performance in certain cases, it struggles with likelihood estimation, as shown in Table 3. Conversely, as discussed in the previous section, NPR-DDPM sometimes achieves superior likelihood results in Table 3, but its image generation quality is significantly worse than OCM-DDPM. Therefore, the proposed OCM-DDPM provides a more balanced model, offering a better trade-off between generation quality and likelihood estimation.
Table 3: The NLL (bits/dim) $\downarrow$ across various datasets using different sampling steps.
CIFAR10 (LS) CIFAR10 (CS)
# timesteps $K$ 10 25 50 100 200 1000 10 25 50 100 200 1000
DDPM, $\tilde{\beta}$ 74.95 24.98 12.01 7.08 5.03 3.73 75.96 24.94 11.96 7.04 4.95 3.60
DDPM, $\beta$ 6.99 6.11 5.44 4.86 4.39 3.75 6.51 5.55 4.92 4.41 4.03 3.54
A-DDPM 5.47 4.79 4.38 4.07 3.84 3.59 5.08 4.45 4.09 3.83 3.64 3.42
NPR-DDPM 5.40 4.64 4.25 3.98 3.79 3.57 5.03 4.33 3.99 3.76 3.59 3.41
SN-DDPM 30.79 11.83 7.13 5.24 4.39 3.74 90.85 19.81 9.72 6.72 5.58 4.73
OCM-DDPM 5.32 4.63 4.25 3.97 3.78 3.57 4.99 4.34 3.99 3.76 3.59 3.41
CelebA 64x64 ImageNet 64x64
# timesteps $K$ 10 25 50 100 200 1000 25 50 100 200 400 4000
DDPM, $\tilde{\beta}$ 33.42 13.09 7.14 4.60 3.45 2.71 105.87 46.25 22.02 12.10 7.59 3.89
DDPM, $\beta$ 6.67 5.72 4.98 4.31 3.74 2.93 5.81 5.20 4.70 4.31 4.04 3.65
A-DDPM 4.54 3.89 3.48 3.16 2.92 2.66 4.78 4.42 4.15 3.95 3.81 3.61
NPR-DDPM 4.46 3.78 3.40 3.11 2.89 2.65 4.66 4.22 3.96 3.80 3.71 3.60
SN-DDPM 18.09 8.05 5.29 4.05 3.40 2.84 4.56 4.18 3.95 3.80 3.71 3.63
OCM-DDPM 4.69 3.86 3.43 3.13 2.90 2.66 4.45 4.15 3.93 3.79 3.70 3.59
5.3 Image Modelling with Latent Diffusion Models
In this section, we apply our methods to latent diffusion models (Vahdat et al., 2021; Rombach et al., 2022) and with ImageNet $256{\times}256$ to demonstrate the scalability of our method. Specifically, we compare our methods to other approaches within the DiT architecture (Peebles & Xie, 2023), evaluating sample quality using the FID score and diversity using the Recall score (Sajjadi et al., 2018). We focus on conditional generation with classifier-free guidance (CFG), which is consistent to Peebles & Xie (2023).
It is important to note that DiT was initially trained using the Improved DDPM (I-DDPM) algorithm (Nichol & Dhariwal, 2021). Therefore, the results of I-DDPM reflect the original performance of the pre-trained DiT model [^1] We use the DiT XL/2 checkpoint released in https://github.com/facebookresearch/DiT. provided by Peebles & Xie (2023). For DDPM and DDIM sampling, we discard the learned covariance in DiT and use only the learned noise prediction neural network for the posterior mean. In contrast, our OCM-DPM retains the same mean function while learning the covariance using the optimal covariance matching objective. We then generate samples with 10 sampling steps and report FID and Recall scores across different CFG coefficients on the ImageNet 256x256 dataset. The results are displayed in Figure 4, with DDPM-$\beta$ excluded due to its poor performance. It shows that our OCM-DPM method achieves the best FID performance with $\text{CFG}=2.0$. Although DDPM-$\tilde{\beta}$ and DDIM perform better in terms of FID at $\text{CFG}=1.5$, their Recall scores are lower than those of the proposed OCM methods. This discrepancy arises partly because DDPM and DDIM model
Figure 4: Results of DiT training on ImageNet 256x256. We generate samples using 10 timesteps with varying CFG coefficients (see Table 11 for exact numerical values).
the inverse process with deterministic variance, whereas OCM methods explicitly learn the variance from data, allowing for more accurate density estimation and generating more diverse results. As Peebles & Xie (2023) report that DiT achieves the best FID performance with $\text{CFG}=1.5$, we further evaluate performance with different sampling steps under $\text{CFG}=1.5$, as shown in Table 12. The results indicate that our OCM-DPM methods offer a great improvement in sample quality and diversity.
Table 4: The least number of timesteps $\downarrow$ required to achieve an FID around 6 (along with the corresponding FID). To account for the additional time cost incurred by the covariance prediction network, we multiply the results by the ratio of the time cost per single timestep, reflecting the extra computational overhead (see Section B.1 for details).
Method CIFAR10 CelebA 64x64 LSUN Bedroom ImageNet 256x256
DDPM 90 (6.12) $>$ 200 130 (6.06) 21 (5.89)
DDIM 30 (5.85) $>$ 100 Best FID > 6 11 (5.58)
Improved DDPM 45 (5.96) Missing model 90 (6.02) 22 (6.08)
Analytic-DPM 25 (5.81) 55 (5.98) 100 (6.05) Missing model
NPR-DPM 1.002$\times$23 (5.76) 1.013$\times$50 (6.04) 1.021$\times$90 (6.01) Missing model
SN-DPM 1.005$\times$17 (5.81) 1.019$\times$22 (5.96) 1.114$\times$92 (6.02) Missing model
OCM-DPM (ours) 1.003$\times$16 (5.83) 1.015$\times$21 (5.94) 1.112$\times$90 (6.04) 1.007$\times$10 (5.33)
To provide an overview of the superiority of our method, we compare the minimum sampling step to achieve the FID score (Heusel et al., 2017) close to $6$ across different approaches (see Section B.1 for details). As shown in Table 4, our method requires the fewest denoising steps in most settings. Moreover, we emphasize that compared to recent advanced methods like SN-DPM and NPR-DPM, our approach is the only one that consistently delivers both competitive likelihood and FID.
Table 5: FID score $\downarrow$ across various datasets using different sampling steps.
CIFAR10 (LS) CIFAR10 (CS)
# timesteps $K$ 10 25 50 100 200 1000 10 25 50 100 200 1000
DDPM, $\tilde{\beta}$ 44.45 21.83 15.21 10.94 8.23 5.11 34.76 16.18 11.11 8.38 6.66 4.92
DDPM, $\beta$ 233.41 125.05 66.28 31.36 12.96 3.04 205.31 84.71 37.35 14.81 5.74 3.34
A-DDPM 34.26 11.60 7.25 5.40 4.01 4.03 22.94 8.50 5.50 4.45 4.04 4.31
NPR-DDPM 32.35 10.55 6.18 4.52 3.57 4.10 19.94 7.99 5.31 4.52 4.10 4.27
SN-DDPM 24.06 6.91 4.63 3.67 3.31 3.65 16.33 6.05 4.17 3.83 3.72 4.07
OCM-DDPM 24.94 9.19 5.95 4.36 3.48 3.98 14.32 5.54 4.10 3.84 3.75 4.18
DDIM 21.31 10.70 7.74 6.08 5.07 4.13 34.34 16.68 10.48 7.94 6.69 4.89
A-DDIM 14.00 5.81 4.04 3.55 3.39 3.74 26.43 9.96 6.02 4.88 4.92 4.66
NPR-DDIM 13.34 5.38 3.95 3.53 3.42 3.72 22.81 9.47 6.04 5.02 5.06 4.62
SN-DDIM 12.19 4.28 3.39 3.23 3.22 3.65 17.90 7.36 5.16 4.63 4.63 4.51
OCM-DDIM 10.66 4.35 3.48 3.27 3.29 3.74 16.70 6.71 4.72 4.30 4.54 4.53
CelebA 64x64 ImageNet 64x64
# timesteps $K$ 10 25 50 100 200 1000 25 50 100 200 400 4000
DDPM, $\tilde{\beta}$ 36.69 24.46 18.96 14.31 10.48 5.95 29.21 21.71 19.12 17.81 17.48 16.55
DDPM, $\beta$ 294.79 115.69 53.39 25.65 9.72 3.16 170.28 83.86 45.04 28.39 21.38 16.38
A-DDPM 28.99 16.01 11.23 8.08 6.51 5.21 32.56 22.45 18.80 17.16 16.40 16.34
NPR-DDPM 28.37 15.74 10.89 8.23 7.03 5.33 28.27 20.89 18.06 16.96 16.32 16.38
SN-DDPM 20.60 12.00 7.88 5.89 5.02 4.42 27.58 20.74 18.04 16.61 16.37 16.22
OCM-DDPM 21.55 12.71 9.24 6.97 5.92 5.04 28.02 20.81 17.98 16.74 16.32 16.31
DDIM 20.54 13.45 9.33 6.60 4.96 3.40 26.06 20.10 18.09 17.84 17.74 19.00
A-DDIM 15.62 9.22 6.13 4.29 3.46 3.13 25.98 19.23 17.73 17.49 17.44 18.98
NPR-DDIM 14.98 8.93 6.04 4.27 3.59 3.15 28.84 19.62 17.63 17.42 17.30 18.91
SN-DDIM 10.20 5.48 3.83 3.04 2.85 2.90 28.07 19.38 17.53 17.23 17.23 18.89
OCM-DDIM 10.28 5.72 4.42 3.54 3.17 3.03 28.28 19.62 17.71 17.42 17.26 19.02
6 Related Work and Future Directions
In addition to the covariance estimation methods discussed in Section 4, there are other approaches to accelerate sampling in diffusion models. One approach involves using faster numerical solvers for differential equations with continuous timesteps (Jolicoeur-Martineau et al., 2021; Liu et al., 2022; Lu et al., 2022). Another strategy, inspired by Schrödinger bridge (Wang et al., 2021; De Bortoli et al., 2021), is to introduce a nonlinear, trainable forward diffusion process. Additionally, combining diffusion models with GANs enhances GAN sample quality while speeding up diffusion model sampling (Xiao et al., 2021; Wang et al., 2022).
Recently, distillation techniques have gained popularity, achieving state-of-the-art in one-step generation (Zhou et al., 2024). There are two prominent types of distillation methods. The first is trajectory distillation (Salimans & Ho, 2022; Berthelot et al., 2023; Song et al., 2023; Heek et al., 2024; Kim et al., 2023; Li & He, 2024), which focuses on accelerating the process of solving differential equations. The second type involves distillation techniques that utilize a one-step implicit latent variable model as the student model (Luo et al., 2024; Salimans et al., 2024; Xie et al., 2024; Zhou et al., 2024) and distills the diffusion process into the student model by minimizing the spread divergence family (Zhang et al., 2020; 2019) through score estimation (Poole et al., 2022; Wang et al., 2024).
Although these distillation methods typically offer faster generation speeds (fewer NFEs) compared to covariance estimation methods, they often lack tractable density or likelihood estimation. This presents a challenge for generative modeling applications where likelihood or density estimation is crucial. For example, in AI4Science problems involving energy-based data, accurate model density estimation is essential for enabling importance sampling to correct sample bias, as discussed in (Zhang et al., 2024a; Vargas et al., 2023; Chen et al., 2024; Akhound-Sadegh et al., 2024). Another example is diffusion model-based data compression, where better likelihood corresponds to better compression rates (Townsend et al., 2019; Zhang et al., 2021; Ho et al., 2020; Kingma et al., 2021). In these cases, the proposed OCM-DDPM can be used to provide better likelihood estimates, resulting in improved task performance. Additionally, a recent paper (Salimans et al., 2024) shows how moment-matching improves distillation, achieving state-of-the-art diffusion quality with first-order matching. Our method could extend this to higher-order moment matching, potentially accelerating distillation training. We leave it as a promising direction for future work.
Beyond image modelling, our method can be straightforwardly applied to accelerate large-scale video diffusion models (Blattmann et al., 2023; Chen et al., 2023), which are based on latent diffusion models. Furthermore, a recent study (Zhao et al., 2024) shows that covariance estimation is crucial in addressing the image-leakage issue in image-to-video generation problems, which presents another important application of our method that we will explore in future work.
7 Conclusion
In this paper, we proposed a novel method for learning the optimal diagonal covariance of the denoising distribution, which offers improved accuracy compared to other covariance learning approaches. Our results demonstrate that enhanced covariance estimation leads to better generation quality and diversity, all while reducing the number of required generation steps. We validated the scalability of our method across different problem domains and scales. Additionally, we discussed its connection to various acceleration techniques and highlighted several promising application areas for future exploration.
Acknowldgement
We want to thank Wenlin Chen for the useful discussions. MZ and DB acknowledge funding from the Cisco Centre of Excellence. ZO is supported by the Lee Family Scholarship.
Appendix for “Improved Diffusion Models with Optimal Covariance Matching"
[…]
Appendix A Abstract Proof and Derivations
A.1 Derivations of the Analytical Full Covariance Identity
See 1
Proof.
This proof generalizes the original analytical covariance identity proof discussed in Zhang et al. (2024b). Using the fact that $\nabla_{\tilde{x}}q(\tilde{x}|x)=-\frac{1}{\sigma^{2}}(\tilde{x}-\alpha x)q(\tilde{x}|x)$ and the Tweedie’s formula $\nabla_{\tilde{x}}\log q(\tilde{x})=\frac{1}{\sigma^{2}}\left(\alpha\mathbb{E}_{q(x|\tilde{x})}[x]-\tilde{x}\right)$, we can expand the hessian of the $\log q_{\theta}(\tilde{x})$:
$\nabla^{2}_{\tilde{x}}\log q(\tilde{x})$ $=-\frac{1}{\sigma^{2}}\int\nabla_{\tilde{x}}\left((\tilde{x}-\alpha x)\frac{q(\tilde{x}|x)q(x)}{q(\tilde{x})}\right)\mathop{}\!\mathrm{d}{x}$
$=-\frac{1}{\sigma^{2}}\int\frac{q(\tilde{x}|x)q(x)}{q(\tilde{x})}\mathop{}\!\mathrm{d}{x}$ $-\frac{1}{\sigma^{2}}\int(\tilde{x}-\alpha x)\frac{\nabla_{\tilde{x}}q(\tilde{x}|x)q(\tilde{x})q(x)-\nabla_{\tilde{x}}q(\tilde{x})q(\tilde{x}|x)q(x)}{q^{2}(\tilde{x})}\mathop{}\!\mathrm{d}{x}$
$\Longrightarrow\sigma^{2}\nabla_{\tilde{x}}^{2}\log q(\tilde{x})+1$ $=-\int(\tilde{x}-\alpha x)\frac{\nabla_{\tilde{x}}q(\tilde{x}|x)q(x)-\nabla_{\tilde{x}}\log q(\tilde{x})q(\tilde{x}|x)q(x)}{q(\tilde{x})}\mathop{}\!\mathrm{d}{x}$
$=-\int(\tilde{x}-\alpha x)$ $\frac{-\frac{1}{\sigma^{2}}(\tilde{x}-\alpha x)q(\tilde{x}|x)q(x)+\frac{1}{\sigma^{2}}(\tilde{x}-\alpha\mathbb{E}_{q(x|\tilde{x})}[x])q(\tilde{x}|x)q(x)}{q(\tilde{x})}\mathop{}\!\mathrm{d}{x}$
$\Longrightarrow\sigma^{4}\nabla_{\tilde{x}}^{2}\log q(\tilde{x})+\sigma^{2}I$ $=\int\left((\tilde{x}-\alpha x)^{2}-(\tilde{x}-\alpha x)(\tilde{x}-\alpha\mathbb{E}_{q(x|\tilde{x})}[x])\right)q(x|\tilde{x})\mathop{}\!\mathrm{d}{x}$
$=\alpha^{2}\mathbb{E}_{q(x|\tilde{x})}[x^{2}]-\alpha^{2}\mathbb{E}_{q(x|\tilde{x})}[x]^{2}\equiv\alpha^{2}\Sigma(\tilde{x})$
Therefore, we obtain the analytical full covariance identity: $\Sigma_{q}(\tilde{x})=\left(\sigma^{4}\nabla^{2}_{\tilde{x}}\log q(\tilde{x})+\sigma^{2}I\right)/\alpha^{2}.$
A.2 Vadality of the OCM Objective
See 2
Proof.
Recall that to learn the optimal covariance, we can minimize the following loss function
$\min_{\phi}\mathbb{E}_{q(\tilde{x})}\left\lVert h_{\phi}(\tilde{x})-\mathbb{E}_{q(v)}[v\odot H(\tilde{x})v]\right\rVert_{2}^{2}$ (25)
We call this the grounded objective because it remains unbiased and consistent, i.e., $h_{\phi^{*}}=\text{diag}(H(\tilde{x}))$, when the inner expectation is approximated with infinite Monte Carlo samples. Using Jensen’s inequality, we can show that the OCM objective defined in (11) provides an upper bound for the grounded objective
$\mathbb{E}_{q(\tilde{x})}\left\lVert h_{\phi}(\tilde{x})-\mathbb{E}_{q(v)}[v\odot H(\tilde{x})v]\right\rVert_{2}^{2}$ $=\mathbb{E}_{q(\tilde{x})}\left\lVert\mathbb{E}_{q(v)}[h_{\phi}(\tilde{x})-v\odot H(\tilde{x})v]\right\rVert_{2}^{2}$
$\leq\mathbb{E}_{q(\tilde{x})}\mathbb{E}_{q(v)}\left\lVert h_{\phi}(\tilde{x})-v\odot H(\tilde{x})v\right\rVert_{2}^{2}$
$=L_{\text{ocm}}(\phi).$
Thus, minimizing the OCM objective also minimizes the grounded objective, leading to a more accurate approximation of the diagonal Hessian. We then show that these two objectives are equivalent when attaining their optimal. To see this, we can expand the OCM objective
$L_{\text{ocm}}(\phi)$ $=\mathbb{E}_{p(v)q(\tilde{x})}||h_{\phi}(\tilde{x})-v\odot H(\tilde{x})v||_{2}^{2}$
$=\mathbb{E}_{q(\tilde{x})}||h_{\phi}(\tilde{x})||_{2}^{2}-2\mathbb{E}_{p(v)q(\tilde{x})}[h_{\phi}(\tilde{x})^{T}(v\odot H(\tilde{x})v)]+c$
$=\mathbb{E}_{q(\tilde{x})}||h_{\phi}(\tilde{x})||_{2}^{2}-2\mathbb{E}_{q(\tilde{x})}[h_{\phi}(\tilde{x})^{T}\text{diag}(H(\tilde{x}))]+c$
$=\mathbb{E}_{q(\tilde{x})}||h_{\phi}(\tilde{x})-\text{diag}(H(\tilde{x}))||_{2}^{2}+c^{\prime},$
where $c,c^{\prime}$ are constants w.r.t. the parameter $\phi$, and line 2 to line 3 follows from the fact that $\mathbb{E}_{p(v)}[v\odot H(x)v]=\text{diag}(H(x))$. Therefore, $L_{\text{ocm}}(\phi)$ attains optimal when $h_{\phi}(\tilde{x})=\text{diag}(H(\tilde{x}))$ for all $\tilde{x}\sim q(\tilde{x})$. ∎
A.3 Connection to SN-DDPM
In this section, we showcase the connection between OCM-DDPM and SN-DDPM through the second-order Tweedie’s formula. Before delving into it, we present the following lemmas, which are essential for our derivation.
Lemma 1 (First order Tweedie’s formula (Efron, 2011)).
Let $q(\tilde{x}|x)=\mathcal{N}(\tilde{x}|\sqrt{\alpha}x,\beta I)$, we have the mean of the inverse density equals
$\mathbb{E}_{q(x|\tilde{x})}[x]=\frac{1}{\sqrt{\alpha}}(\tilde{x}+\beta\nabla_{\tilde{x}}\log q(\tilde{x})).$ (26)
Lemma 2 (Second order Tweedie’s formula).
Let $q(\tilde{x}|x)=\mathcal{N}(\tilde{x}|\sqrt{\alpha}x,\beta I)$, we have the second moment of the inverse density equals
$\mathbb{E}_{q(x|\tilde{x})}[xx^{T}]=\frac{1}{\alpha}\left(\tilde{x}\tilde{x}^{T}+\beta s_{1}(\tilde{x})\tilde{x}^{T}+\beta\tilde{x}s_{1}(\tilde{x})^{T}+\beta^{2}s_{2}(\tilde{x})+\beta^{2}s_{1}(\tilde{x})s_{1}(\tilde{x})^{T}+\beta I\right),$ (27)
where $s_{1}(\tilde{x})\equiv\nabla_{\tilde{x}}\log q(\tilde{x})$ and $s_{2}(\tilde{x})\equiv\nabla_{\tilde{x}}^{2}\log q(\tilde{x})$.
Proof.
The proof follows that in (Meng et al., 2021, Appendix B) with the generalization to the scaled Gaussian convolutions. Specifically, we first reparametrized $q(\tilde{x}|x)$ as a exponential distribution
$q(\tilde{x}|\eta)=e^{\eta^{T}\tilde{x}-\psi(\eta)}q_{0}(\tilde{x}),$ (28)
where $\eta=\frac{\sqrt{\alpha}}{\beta}x,q_{0}(\tilde{x})\propto e^{-\frac{1}{2\beta}\tilde{x}^{T}\tilde{x}}$, and $\psi(\eta)$ denotes the partition function. By applying the Bayes rule $q(\eta|\tilde{x})=\frac{q(\tilde{x}|\eta)p(\eta)}{q(\tilde{x})}$, we have the corresponding posterior
$q(\eta|\tilde{x})=e^{\eta^{T}\tilde{x}-\psi(\eta)-\lambda(\tilde{x})}p(\eta),$ (29)
where $\lambda(\tilde{x})\equiv\log q(\tilde{x})-\log q_{0}(\tilde{x})$. Since $\int q(\eta|\tilde{x})\mathrm{d}\eta=1$, by taking the derivative w.r.t. $\tilde{x}$ on both sides, we have
$\int\left(\eta-\nabla_{\tilde{x}}\lambda(\tilde{x})\right)^{T}q(\eta|\tilde{x})=0,$ (30)
which implies that $\mathbb{E}[\eta|\tilde{x}]=\nabla_{\tilde{x}}\lambda(\tilde{x})$. Taking the derivative w.r.t. $\tilde{x}$ on both sides again, we have
$\int\eta\left(\eta-\nabla_{\tilde{x}}\lambda(\tilde{x})\right)^{T}q(\eta|\tilde{x})=\nabla_{\tilde{x}}^{2}\lambda(\tilde{x}),$ (31)
which implies that $\mathbb{E}[\eta\eta^{T}|\tilde{x}]=\nabla_{\tilde{x}}^{2}\lambda(\tilde{x})+\nabla_{\tilde{x}}\lambda(\tilde{x})\nabla_{\tilde{x}}^{T}\lambda(\tilde{x})$. By substituting $\eta=\frac{\sqrt{\alpha}}{\beta}x,\nabla_{\tilde{x}}\lambda(\tilde{x})=s_{1}(\tilde{x})+\frac{1}{\beta}\tilde{x}$, and $\nabla_{\tilde{x}}^{2}\lambda(\tilde{x})=s_{2}(\tilde{x})+\frac{1}{\beta}I$, we get the result as desired. ∎
Lemma 3 (Convert the covariance of $q(\tilde{x}|x)$ to the hessian of $q(\tilde{x})$).
Let $q(\tilde{x}|x)=\mathcal{N}(\tilde{x}|\sqrt{\alpha}x,\beta I)$, we have the covariance of the inverse density equals
$\mathrm{Cov}_{q(x|\tilde{x})}[x]=\frac{\beta}{\alpha}(I+\beta\nabla_{\tilde{x}}^{2}\log q(\tilde{x})).$ (32)
Proof.
Let $s_{1}(\tilde{x})\equiv\nabla_{\tilde{x}}\log q(\tilde{x})$ and $s_{2}(\tilde{x})\equiv\nabla_{\tilde{x}}^{2}\log q(\tilde{x})$. We have
$\mathrm{Cov}_{q(x|\tilde{x})}[x]$ $=\frac{\beta^{2}}{\alpha}\mathrm{Cov}_{q(x|\tilde{x})}\left[\frac{\tilde{x}-\sqrt{\alpha}x}{\beta}\right]$
$=\frac{\beta^{2}}{\alpha}\left(\mathbb{E}_{q(x|\tilde{x})}\left(\frac{\tilde{x}\!-\!\sqrt{\alpha}x}{\beta}\right)\left(\frac{\tilde{x}\!-\!\sqrt{\alpha}x}{\beta}\right)^{\!T}\!\!\!\!-\mathbb{E}_{q(x|\tilde{x})}\left(\frac{\tilde{x}\!-\!\sqrt{\alpha}x}{\beta}\right)\mathbb{E}_{q(x|\tilde{x})}\left(\frac{\tilde{x}\!-\!\sqrt{\alpha}x}{\beta}\right)^{\!T}\right)$
$=\frac{\beta^{2}}{\alpha}\left(\frac{1}{\beta^{2}}\mathbb{E}_{q(x|\tilde{x})}\left(\tilde{x}-\sqrt{\alpha}x\right)\left(\tilde{x}-\sqrt{\alpha}x\right)^{T}-s_{1}(\tilde{x})s_{1}(\tilde{x})^{T}\right)$
$=\frac{\beta^{2}}{\alpha}\left(\frac{1}{\beta^{2}}\left(\tilde{x}\tilde{x}^{T}-2\tilde{x}(\tilde{x}+\beta s_{1}(\tilde{x}))^{T}+\alpha\mathbb{E}_{q(x|\tilde{x})}[xx^{T}]\right)-s_{1}(\tilde{x})s_{1}(\tilde{x})^{T}\right)$
$=\frac{\beta^{2}}{\alpha}\left(\frac{1}{\beta^{2}}\left(\beta^{2}s_{2}(\tilde{x})+\beta^{2}s_{1}(\tilde{x})s_{1}(\tilde{x})^{T}+\beta I\right)-s_{1}(\tilde{x})s_{1}(\tilde{x})^{T}\right)$
$=\frac{\beta^{2}}{\alpha}\left(\frac{1}{\beta}I+s_{2}(\tilde{x})\right)\equiv\frac{\beta}{\alpha}\left(I+\beta\nabla_{\tilde{x}}^{2}\log q(\tilde{x})\right),$
where line 2 to line 3 follows Lemma 1 and line 4 to line 5 follows Lemma 2. ∎
It is noteworthy that line 3 in the proof of Lemma 3 also showcases the connection between the covariance of $q(\tilde{x}|x)$ and the score of $q(\tilde{x})$:
$\mathrm{Cov}_{q(x|\tilde{x})}[x]=\frac{\beta^{2}}{\alpha}\left(\frac{1}{\beta}\mathbb{E}_{q(x|\tilde{x})}[\epsilon\epsilon^{T}]-\nabla_{\tilde{x}}\log q(\tilde{x})\nabla_{\tilde{x}}\log q(\tilde{x})^{T}\right),$ (33)
where $\epsilon=(\tilde{x}-\sqrt{\alpha}x)/\sqrt{\beta}$. Now we can apply Equations 32 and 33 to establish the connection between OCM-DDPM and SN-DDPM, as demonstrated in the following theorem.
Theorem 3 (Connection between OCM-DDPM and SN-DDPM).
Suppose $q(x_{0:T})$ is defined as Equation 1, and the pre-trained score function is well-learned, i.e., $\nabla_{x_{t}}\log p_{\theta}(x_{t})=\nabla_{x_{t}}\log q(x_{t}),\forall x_{t}$. Let $h_{\phi},g_{\psi}$ be parameterized neural networks. OCM-DDPM learns $h_{\phi}$ by minimizing the objective $\mathcal{L}_{\text{OCM}}(\phi)$ as in Equation 13, and SN-DDPM learns $g_{\psi}$ by minimizing the objective $\mathcal{L}_{\text{SN}}(\psi)$ as in Equation 23. Then in optimal training with $\phi^{*}=\operatorname*{argmin}\mathcal{L}_{\text{OCM}}(\phi)$ and $\psi^{*}=\operatorname*{argmin}\mathcal{L}_{\text{SN}}(\psi)$, we have the optimal diagonal covariance of OCM-DDPM $\Sigma_{t-1}(x_{t};\phi^{*})=\frac{(1-\alpha_{t})^{2}h_{\phi^{*}}(x_{t})+(1-\alpha_{t})I}{{\alpha}_{t}}$ and that of SN-DDPM $\Sigma_{t-1}(x_{t};\psi^{*})=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}+\frac{\beta_{t}^{2}}{\alpha_{t}}\!\left(\!\frac{g_{\psi^{*}}(x_{t})}{1\!-\!\bar{\alpha}_{t}}\!-\!\nabla_{x_{t}}\log\!p_{\theta}(x_{t})^{2}\!\right)$ are identical.
Proof.
In optimal training, we know that $h_{\phi^{*}}(x_{t})=\mathrm{diag}(\nabla_{x_{t}}^{2}\log q(x_{t}))$ and $g_{\psi^{*}}(x_{t})=\mathbb{E}_{q(x_{0}|x_{t})}[\epsilon_{t}^{2}]$. Bao et al. (2022b) show that the covariance of the denoising density $q(x_{t-1}|x_{t})$ has a closed form (see Lemma 13 in Bao et al. (2022b))
$\mathrm{Cov}_{q(x_{t-1}|x_{t})}[x_{t-1}]=\lambda_{t}^{2}I+\gamma_{t}^{2}\mathrm{Cov}_{q(x_{0}|x_{t})}[x_{0}],$ (34)
where $\lambda_{t}^{2}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}$ and $\gamma_{t}=\sqrt{\bar{\alpha}_{t-1}}-\sqrt{1-\bar{\alpha}_{t-1}-\lambda_{t}^{2}}\sqrt{\frac{\bar{\alpha}_{t}}{1-\bar{\alpha}_{t}}}=\sqrt{\bar{\alpha}_{t-1}}\frac{\beta_{t}}{1-\bar{\alpha}_{t}}$. Since $q(x_{t}|x_{0})=\mathcal{N}(\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})I)$, applying Lemma 3 gives
$\mathrm{diag}(\mathrm{Cov}_{q(x_{0}|x_{t})}[x_{0}])=\frac{(1-\bar{\alpha}_{t})}{\bar{\alpha}_{t}}(I+(1-\bar{\alpha}_{t})h_{\phi^{*}}(x_{t})).$
Substituting it into Equation 34, we have
$\mathrm{diag}(\mathrm{Cov}_{q(x_{t-1}|x_{t})}[x_{t-1}])$ $=\frac{(1-\alpha_{t})^{2}h_{\phi^{*}}(x_{t})+(1-\alpha_{t})I}{{\alpha}_{t}}=\Sigma_{t-1}(x_{t};\phi^{*})$
as desired. Alternatively, applying Equation 33, we have
$\mathrm{diag}(\mathrm{Cov}_{q(x_{0}|x_{t})}[x_{0}])=\frac{(1-\bar{\alpha}_{t})^{2}}{\bar{\alpha}_{t}}\left(\frac{g_{\psi^{*}}(x_{t})}{1-\bar{\alpha}_{t}}-\nabla_{x_{t}}\log\!p_{\theta}(x_{t})^{2}\right).$
Substituting it into Equation 34 gives
$\mathrm{diag}(\mathrm{Cov}_{q(x_{t-1}|x_{t})}[x_{t-1}])=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}+\frac{\beta_{t}^{2}}{\alpha_{t}}\!\left(\!\frac{g_{\psi^{*}}(x_{t})}{1\!-\!\bar{\alpha}_{t}}\!-\!\nabla_{x_{t}}\log\!p_{\theta}(x_{t})^{2}\!\right)=\Sigma_{t-1}(x_{t};\psi^{*}).$
Therefore, $\mathrm{diag}(\mathrm{Cov}_{q(x_{t-1}|x_{t})}[x_{t-1}])=\Sigma_{t-1}(x_{t};\phi^{*})=\Sigma_{t-1}(x_{t};\psi^{*})$ as desired. ∎
Remark.
Theorem 3 establishes the connection between OCM-DDPM and SN-DDPM using the second-order Tweedie’s formula. This connection allows OCM-DDPM to utilize the same covariance clipping trick as in SN-DDPM (see Section B.2 for details). Additionally, as highlighted in Bao et al. (2022a), SN-DDPM suffers from error amplification in the quadratic term, a limitation not present in OCM-DDPM. As shown in Figures 3 and 3, OCM-DDPM demonstrates superior covariance estimation accuracy and likelihood performance compared to SN-DDPM, underscoring the advantages of the proposed optimal covariance matching objective.
Appendix B Details of Experiments
B.1 Details of Model Architectures
In this section, we describe our model architectures in detail. Following the approach in Bao et al. (2022a), our model comprises a pretrained score neural network with fixed parameters, along with a trainable diagonal Hessian prediction network built on top of it.
Table 6: Source of pretrained score prediction networks used in our experiments.
Provided By
CIFAR10 (LS) Bao et al. (2022b)
CIFAR10 (CS) Bao et al. (2022b)
CelebA 64X64 Song et al. (2020)
ImageNet 64X64 Nichol & Dhariwal (2021)
LSUN Bedroom Ho et al. (2020)
ImageNet 256x156 Peebles & Xie (2023)
Details of Pretrained Score Prediction Networks. Table 6 lists the pretrained neural networks utilized in our experiments. These models parameterize the noise prediction $\epsilon_{\theta}(x)$, allowing us to derive the score prediction as $s_{\theta}(x_{t})=\nabla_{x}\log p_{\theta}(x_{t})=-\frac{\epsilon_{\theta}(x_{t})}{\sqrt{1-\bar{\alpha}_{t}}}$, following the forward process defined in Equation 1. It is important to note that the pretrained networks for ImageNet 64x64 and 256x256 include both noise prediction networks and covariance networks. In our model architectures, we utilize only the noise prediction networks.
Table 7: Architecture details of our models.
$\mathrm{NN}_{1}$ $\mathrm{NN}_{2}$
CIFAR10 (LS) $\mathrm{Conv}$ $\mathrm{Conv}$
CIFAR10 (CS) $\mathrm{Conv}$ $\mathrm{Conv}$
CelebA 64X64 $\mathrm{Conv}$ $\mathrm{Conv}$
ImageNet 64X64 $\mathrm{Conv}$ $\mathrm{Res}$+$\mathrm{Conv}$
LSUN Bedroom $\mathrm{Conv}$ $\mathrm{Res}$+$\mathrm{Conv}$
ImageNet 256x156 $\mathrm{AdaLN}$+$\mathrm{Linear}$ $\mathrm{AdaLN}$+$\mathrm{Linear}$
Details of Diagonal Hessian Prediction Networks. For fair comparisons, we follow the parameterization as per Bao et al. (2022a) for all models excluding the one on ImageNet 256x256, which was not explored in their paper. The architecture details of $\mathrm{NN}_{1}$ and $\mathrm{NN}_{2}$ are provided in Table 7, where $\mathrm{Conv}$ denotes the convolutional layer, $\mathrm{Res}$ denotes the residual block (He et al., 2016), $\mathrm{AdaLN}$ denotes the adaptive layer norm block (Peebles & Xie, 2023), and $\mathrm{Linear}$ denotes the linear layer.
Table 8: The number of parameters and the averaged time (ms) to run a model function evaluation. All are evaluated with a batch size of 64 on one A100-80GB GPU.
Score prediction network Score & SN prediction networks Score & Diagonal Hessian Prediction networks
CIFAR10 (LS) 52.54 M / 44.37 ms 52.55 M / 44.61 ms (+0.5%) 52.55 M / 44.51 ms (+0.3%)
CIFAR10 (CS) 52.54 M / 45.13 ms 52.55 M / 45.24 ms (+0.2%) 52.55 M / 45.23 ms (+0.2%)
CelebA 64x64 78.70 M / 67.88 ms 78.71 M / 69.15 ms (+1.9%) 78.71 M / 68.89 ms (+1.5%)
ImageNet 64x64 121.06 M / 106.58 ms 121.49 M / 112.53 ms (+5.6%) 121.49 M / 112.23 ms (+5.3%)
LSUN Bedroom 113.67 M / 692.58 ms 114.04 M / 771.55 ms (+11.4%) 114.04 M / 770.73 ms (+11.2%)
ImageNet 256x256 675.13 M / 22.84 ms Missing model 677.80 M / 23.01 ms (+0.7%)
Cost of Memory and Inference Time. In Table 8, we present the number of parameters and inference time for various models. It is evident that the additional memory cost of the diagonal Hessian prediction network is minimal compared to the original diffusion models, which only include the score prediction. Regarding the extra inference cost, it is negligible for CIFAR10, CelebA 64x64, and ImageNet 256x256. The additional time is at most $5.3\%$ on ImageNet 64x64, and $11.2\%$ on LSUN Bedroom, but this is offset by the benefit of requiring fewer sampling steps to achieve an FID around 6, as shown in Table 4. Notably, although the model size on ImageNet 256x256 is the largest, it has the shortest inference time because DiT learns the diffusion model in the latent space.
Details of Table 4. In Table 4, the FID results of baselines are taken from Bao et al. (2022a). Notably, the time cost for a single neural function evaluation is identical across DDPM, DDIM, and Analytic-DPM, all of which use a fixed isotropic covariance in the backward Markov process. The ratio of the time cost to the baselines is based on Table 8. For the FID results of our model, OCM-DPM, we report performance using the DDPM forward process for LSUN Bedroom and the DDIM forward process for the other datasets.
B.2 Details of Training, Inference, and Evaluation
Our training and inference recipes largely follow those outlined in Bao et al. (2022a); Peebles & Xie (2023). Below, we provide a detailed description.
Training Details. We use the AdamW optimizer (Loshchilov, 2017) with a learning rate of 0.0001 and train for 500K iterations across all datasets. The batch sizes are set to 64 for LSUN Bedroom, 128 for CIFAR10, CelebA 64x64, and ImageNet 64x64, and 256 for ImageNet 256x256. During training, checkpoints are saved every 10K iterations, and we select the checkpoint with the best FID on 2048 samples generated with full sampling steps [^2] We use 2048 samples instead of 1000 to ensure that the covariance of the Inception features has full rank.. We train our models using one A100-80G GPU for CIFAR10, CelebA 64x64, and ImageNet 64x64; four A100-80G GPUs for LSUN Bedroom; and eight A100-80G GPUs for ImageNet 256x256.
Sampling Details. As per Bao et al. (2022b), covariance clipping is crucial to the performance in diffusion models with unfixed variance in the backward Markovian. Leveraging the connection between OCM-DPM and SN-DPM established in Theorem 3, we can apply the same clipping strategies outlined in Bao et al. (2022a; b). Specifically, we only display the mean of $p(x_{0}|x_{1})$ at the last sampling and clip the covariance $\Sigma_{1}(x_{2})$ of $p(x_{1}|x_{2})$ such that $\lVert\Sigma_{1}(x_{2})\rVert_{\infty}\mathbb{E}|\epsilon|\leq\frac{2}{255}y$, where $\lVert\cdot\rVert_{\infty}$ denotes the infinity norm and $\epsilon$ is the standard Gaussian noise. Following Bao et al. (2022b), we use $y=2$ on CIFAR10 (LS) and CelebA 64x64 under the DDPM forward process, and use $y=1$ for other cases. For all skip-step sampling methods, we employ the even trajectory (Nichol & Dhariwal, 2021) for selecting the subset of sampling steps.
Evaluation Details. The performance is evaluated on the exponential moving average (EMA) model with a rate of $0.9999$. For computing the negative log-likelihood, we follow Ho et al. (2020); Bao et al. (2022a) by discretizing the last sampling step $p(x_{0}|x_{1})$ to obtain the likelihood of discrete image data and report the upper bound of the negative log-likelihood on the entire test set. The FID score is computed on 50K generated samples. Following Nichol & Dhariwal (2021); Bao et al. (2022a); Peebles & Xie (2023), the reference distribution statistics for FID are calculated using the full training set for CIFAR10 and ImageNet, and 50K training samples for CelebA and LSUN Bedroom. Performance results are reported using the same random seed [^3] We use the official code from https://github.com/baofff/Extended-Analytic-DPM. as in Bao et al. (2022a). Additionally, the variance across different random seeds is provided in Section C.2.
Table 9: The corresponding mean and standard deviation of NLL and FID reported in Tables 3 and 5. Note that the standard deviation is reported under the scale of percentage (%).
CIFAR10 (LS) CIFAR10 (CS)
# timesteps $K$ 10 25 50 100 200 1000 10 25 50 100 200 1000
Mean-DDPM (NLL) 5.33 4.63 4.35 3.97 3.78 3.57 4.99 4.34 3.99 3.76 3.59 3.41
Std-DDPM (NLL) % 0.04 0.03 0.03 0.04 0.04 0.01 0.02 0.01 0.03 0.03 0.02 0.01
Mean-DDPM (FID) 25.07 9.30 5.90 4.37 3.54 4.00 14.46 5.49 4.09 3.83 3.81 4.22
Std-DDPM (FID) % 9.05 7.54 3.48 2.77 4.25 2.30 10.40 7.74 2.66 1.32 5.26 3.08
Mean-DDIM (FID) 10.61 4.31 3.48 3.26 3.25 3.75 16.85 6.70 4.73 4.33 4.56 4.59
Std-DDIM (FID) % 5.17 3.55 3.19 1.87 3.97 0.99 13.22 4.23 2.28 2.94 1.61 5.30
CelebA 64x64 ImageNet 64x64
# timesteps $K$ 10 25 50 100 200 1000 25 50 100 200 400 4000
Mean-DDPM (NLL) 4.69 3.86 3.43 3.13 2.90 2.66 4.45 4.15 3.93 3.79 3.70 3.59
Std-DDPM (NLL) % 0.01 0.01 0.01 0.00 0.01 0.00 0.01 0.00 0.01 0.00 0.01 0.01
Mean-DDPM (FID) 21.58 12.65 9.24 6.96 6.00 5.02 28.01 20.85 18.01 16.76 16.35 16.33
Std-DDPM (FID) % 2.24 7.65 2.29 2.32 5.50 1.37 0.77 4.66 6.82 3.28 3.04 3.21
Mean-DDIM (FID) 10.36 5.69 4.37 3.50 3.11 3.03 28.34 19.68 17.84 17.41 17.36 18.91
Std-DDIM (FID) % 7.40 4.04 5.12 4.88 6.46 1.55 6.59 3.77 9.15 4.08 7.88 5.26
Appendix C Additional Experimental Results
In this section, we present additional experimental results for comparison. First, we present an additional likelihood comparison between I-DDPM, NPR-DDPM, and OCM-DDPM in Section C.1 and discuss the performance variance in Section C.2. Next, we provide further experimental results on latent diffusion models in Section C.3 and conduct qualitative studies by showcasing samples generated by our models in Section C.4.
C.1 Additional Likelihood Comparison
Table 10: Upper bound on the negative log-likelihood (bits/dim) on the ImageNet 256x256 dataset.
# timesteps $K$ 25 50 100 200 400 4000
I-DDPM 18.91 8.46 5.27 4.24 3.86 3.57
NPR-DDPM 4.66 4.22 3.96 3.80 3.71 3.60
SN-DDPM 4.56 4.18 3.95 3.80 3.71 3.63
OCM-DDPM 4.45 4.15 3.93 3.79 3.70 3.59
Here, we provide additional likelihood comparisons against I-DDPM (Nichol & Dhariwal, 2021) on the ImageNet 64x64 dataset. As discussed in Section 4, I-DDPM parameterizes the diagonal covariance by interpolating between $\beta$ and ${\beta}$ and learns the covariance by maximizing the variational lower bound. For ease of comparison, we also include the performance of NPR-DDPM and SN-DDOM, which employ an alternative MSE loss as detailed in Equation 23 to learn the covariance. Table 10 presents the results, showing that OCM-DDPM achieves the best likelihood with fewer sampling steps while maintaining performance comparable to I-DDPM when using full sampling steps. This highlights the superiority of the proposed optimal covariance matching objective in learning a diffusion model for data density estimation.
C.2 Mean and Variance of Performance
In Tables 3 and 5, we evaluate our models using the same random seed as Bao et al. (2022a). To minimize the impact of randomness, we report the mean and standard deviation in Table 9 by repeating the evaluation three times with different seeds.
C.3 More Results on Latent Diffusion Models
We report the performance using 10 sampling steps with varying CFG coefficients in Table 11. The results indicate that our methods perform best at $\text{CFG}=2.0$. While DDPM-$\tilde{\beta}$ and DDIM show strong FID scores, their Recall is lower due to fixed variance, suggesting less diversity in the generated samples. In Table 12, we further compare the performance across different sampling steps at $\text{CFG}=1.5$. The results again demonstrate that our methods, which estimate the optimal covariance from the data, produce more diverse samples while maintaining comparable image quality.
C.4 Generated Samples
In this section, we conduct qualitative studies by showcasing the generated samples from our models using different sampling steps $K$. The results are summarized as follows:
  • In Figure 5, we visualize the generated samples using varying numbers of Monte Carlo samples with the Rademacher estimator (refer to Table 1).
  • In Figure 6, we visualize the training data and the generated samples using the minimum number of sampling steps required to achieve an FID of approximately 6 (refer to Table 4).
  • In Figures 7, 8, 9 and 10, we visualize the generated samples of our models using different number of sampling steps on CIFAR10 (LS), CIFAR10 (CS), CelebA 64x64, and ImageNet 64x64, respectively (refer to Table 5).
  • In Figure 11, we visualize the generated samples on ImageNet 256x256, using 10 sampling steps with different CFG coefficients. (refer to Figures 4 and 11).
  • In Figure 12, we visualize the generated samples on ImageNet 256x256, using varying number of sampling steps with CFG set to $1.5$ (refer to Table 12).
Table 11: Results with varying CFG coefficients using 10 sampling steps on ImageNet 256x256.
FID $\downarrow$ Recall (%) $\uparrow$
cfg $=$ 1.5 1.75 2.0 3.0 4.0 1.5 1.75 2.0 3.0 4.0
DDPM, $\tilde{\beta}$ 30.41 19.26 13.53 11.52 14.76 39.74 35.29 32.08 24.93 19.91
DDPM, $\beta$ 210.28 182.78 158.16 89.86 58.16 16.98 22.28 25.13 23.07 19.48
I-DDPM 44.96 20.71 13.01 9.76 13.75 50.32 43.75 41.08 31.77 25.49
OCM-DDPM 30.55 17.57 11.12 9.52 13.74 49.39 45.87 42.33 32.13 24.96
DDIM 9.41 6.54 6.12 10.49 14.18 49.15 45.82 41.15 29.27 23.16
OCM-DDIM 11.30 6.67 5.33 9.20 13.49 54.50 50.32 46.58 34.53 25.67
Table 12: Results with CFG=$1.5$ across different sampling steps on ImageNet 256x256.
FID $\downarrow$ Recall (%) $\uparrow$
# timesteps $K$ 10 25 50 100 200 250 10 25 50 100 200 250
DDPM, $\tilde{\beta}$ 30.41 4.79 3.55 3.09 3.05 2.97 39.74 49.02 51.79 53.54 54.21 54.34
DDPM, $\beta$ 210.28 42.09 9.43 3.63 2.91 2.75 16.98 46.76 51.67 55.01 55.50 55.21
I-DDPM 44.96 9.01 3.70 2.48 2.25 2.74 50.32 54.78 56.45 57.88 58.76 54.69
OCM-DDPM 30.55 4.96 3.21 2.71 2.50 2.75 49.39 53.75 54.68 55.85 54.97 54.75
DDIM 9.41 2.70 2.33 2.25 2.23 2.20 49.15 56.41 57.00 57.83 57.81 57.99
OCM-DDIM 11.30 3.26 2.56 2.28 2.23 2.18 54.50 58.33 58.09 58.80 58.16 58.56
Figure 5: Generated samples with 10 sampling steps using ifferent Rademacher sample numbers on CIFAR10 (CS) (ref: Table 1).
Figure 6: The training data and generated samples of OCM-DPM with minimum steps to achieve an FID around 6. (ref: Table 4)
Figure 7: Generated samples with different sampling steps on CIFAR10 (LS) (ref: Table 5).
Figure 8: Generated samples with different sampling steps on CIFAR10 (CS) (ref Table 5).
Figure 9: Generated samples with different sampling steps on CelebA 64x64 (ref: Table 5).
Figure 10: Generated samples with different sampling steps on ImageNet 64x64 (ref: Table 5).
Figure 11: Generated samples with 10 sampling steps using different CFG coefficients on ImageNet 256x256 (ref: Table 11)
Figure 12: Generated samples with CFG=$1.5$ using different sampling steps on ImageNet 256x256 (ref: Table 12)